# Value Residual Learning For Alleviating Attention Concentration In Transformers

### v1

@@ Abstract
Transformers can capture long-range dependencies using self-attention, allowing tokens to attend to all others directly. However, stacking multiple attention layers leads to attention concentration. One natural way to address this issue is to use cross-layer attention, allowing information from earlier layers to be directly accessible to later layers. However, this approach is computationally expensive. To address this problem, we propose Transformer with residual value (ResFormer) which approximates cross-layer attention through adding a residual connection from the values of the the first layer to all subsequent layers. Based on this method, one variant is the Transformer with single layer value (SVFormer), where all layers share the same value embedding from first layer, reducing the $KV$ cache by nearly 50%. Comprehensive empirical evidence demonstrates that ResFormer mitigates attention concentration problem in deeper layers and enhances representation across most layers, outperforming the vanilla Transformer, DenseFormer, and NeuTRENO in training error as well as downstream tasks. SVFormer trains significantly faster than the vanilla Transformer and performs better than other methods like GQA and CLA, with performance influenced by sequence length and cumulative learning rate.  [^1] Code is available at https://github.com/Zcchill/Value-Residual-Learning.
@@ 1 Introduction
Existing solutions to alleviate the over-smoothing problem in Transformer include adding extra regularizers (Nguyen et al., 2023, Shi et al., 2022) and optimizing the information flow within the model (Pagliardini et al., 2024). During the era of convolutional neural network architectures, Stochastic Depth (Huang et al., 2016) reduces the likelihood of over-smoothing by randomly dropping layers during training and DenseNet (Huang et al., 2017) mitigates the impact of over-smoothing by allowing each layer to directly access the hidden states of all preceding layers. Recently, DenseFormer (Pagliardini et al., 2024) adopts the idea of DenseNet when training Transformer. Additionally, NeuTRENO (Nguyen et al., 2023) alleviates over-smoothing through incorporating the difference between the value vectors of the first layer and the current layer to the attention output.
@@
To alleviate the attention concentration from stacking multiple attention layers, an effective method is to use cross-layer attention on information from earlier layers. Given the high computational cost of cross-layer attention, we propose ResFormer as an approximation. ResFormer achieves a similar effect by applying a residual connection between the value vectors of the current layer and the first layer before the attention operation. Experimental results show that ResFormer mitigates the attention concentration effect in deeper layers, maintaining attention dispersion throughout the sequence, as shown in Fig. 1 (Middle).
@@
Figure 2: Simplified illustration of the vanilla Transformer, NeuTRENO, DenseFormer, ResFormer, and SVFormer, with only three-layer structures and no operations other than attention. $A^{i}$, $V^{i}$, and $H^{i}$ denote the attention matrix, value vectors, and attention outputs at the $i$-th layer, respectively. $\oplus$, $\ominus$, and $\otimes$ represent standard matrix addition, subtraction, and multiplication, respectively.
@@ 2.1 Shortcut Connections For Better Information Flow
Related research indicates that for advanced Transformer architectures, although increasing depth continues to yield performance improvements in language modeling tasks, the gains become less significant with further increases (Petty et al., 2024). Furthermore, Zhou et al. (2021) illustrates that a 32-layer ViT underperforms a 24-layer ViT. Depth-Wise Attention (ElNokrashy et al., 2024) allows each query to access the key and value at the same position from previous layers through an attention-like mechanism before the output layer. DenseFormer (Pagliardini et al., 2024) integrates weighted fusion of outputs from all preceding layers after each layer. To explore why increasing depth in Transformers does not yield expected gains, Wang et al. (2022) finds that self-attention acts as a low-pass filter, smoothing token representations in ViTs. Additionally, Shi et al. (2022) investigates over-smoothing from a graph perspective in BERT-based language modeling tasks. NeuTRENO (Nguyen et al., 2023) adds the difference between the value vectors of the first and current layers to each layer’s attention output and significantly alleviates the over-smoothing problem.
@@ 3.2 Efficient Cross Layer Attention
$\vspace{-5mm}\mathbf{U}_{n}$ $=\mathbf{A}_{n}\mathbf{V}_{n}+\frac{1}{n-1}\sum_{i=1}^{n-1}{\mathbf{V}_{i}}.$ (6)
@@ 3.5 SVFormer: Single-layer Value For Half $KV$ Cache
After ResFormer, a natural idea is whether we can remove the value vectors in each layer and have all layers share the value vectors from the first layer. We call this method SVFormer. Similar to ResFormer, SVFormer still adopts standard attention in the first layer and obtain the attention output $\mathbf{U_{n}}$ for $n$-th layer where $n\geq 2$ through
@@
Compared to previous methods, SVFormer is the first method that decouple value vectors from attention. Its main advantage is that it only requires computing and storing the value vectors for the first layer, saving nearly half of the $KV$ cache during inference. Similar methods like CLA reduce $KV$ cache by sharing both of the key and value vectors every two layers, but the results in Fig. 4 show that sharing values has less negative impact on performance compared with sharing keys.
@@ 4.1.3 Entropy For Analyzing Attention Concentration Effects
where $a_{i}^{{}^{\prime}}={a_{i}}/\Big{(}{\sum_{i=1}^{l}{a_{i}}}\Big{)}$ for $i=1,2,\dots,l$ and the higher the entropy $\bm{e}$, the greater the degree of clustering in $\bm{a}$, i.e., attention matrix $\mathbf{A}$ is more likely to focus on several specific tokens. 1,000 examples with a sequence length of 2,048 in the training dataset are selected for entropy analysis.
@@ 4.1.4 Spectral Decomposition For Analyzing Representations
We trained ResFormer and vanilla Transformer with different model size on data with different sequence lengths. In Fig. 5, ResFormer consistently outperforms vanilla Attention throughout training for different models sizes. Additionally, the results in Fig. 1 (Left) illustrate that ResFormer outperforms DenseFormer and NeuTRENO. Furthermore, integrating ResFormer with NeuTRENO leads to additional performance improvements.
@@ 4.5 Visualization of ResFormer
Figure 9: Left: Distribution of eigenvalues for the value vectors in the first layer of ResFormer and Transformer. Right: Maximum eigenvalue for each layer of ResFormer and Transformer.
To figure out why ResFormer can achieve better performance on language modeling tasks than vanilla Transformer, we conduct visualization based on the eigenvalue decomposition discussed in Section 4.1.4. After sorting the eigenvalues in descending order, we compute the average eigenvalue for each layer across 1,000 randomly sampled pre-train data examples. The results in Fig. 9 indicate that the value vectors generated by most layers of the ResFormer exhibit stronger representational capacity compared to those of the vanilla Transformer. Besides, we analyze the attention concentration effects mentioned in Section 4.1.3 using the same batch of test data. Fig. 1 (Middle) illustrates that the clustering effect of attention increases significantly with the number of layers for the vanilla Transformer, whereas the clustering effect is relatively less pronounced for the ResFormer.
@@ 5 Conclusion
In this paper, we propose the concept of attention concentration, a problem that arises from stacking multiple attention layers. From the perspective of cross-layer attention, we derive ResFormer, which adds a residual connection between the value vectors of the current layer and those of the first layer before the attention operation to alleviate attention concentration. Additionally, we introduce SVFormer, based on ResFormer, which reduces the $KV$ cache by nearly half. We conducted comprehensive experiments on the language modeling task to validate the advantages of these two Transformer variants in different scenarios.
@@ A.1 Token Similarity Analysis
Figure 15: (Left) The average token similarity of hidden states across layers in ResFormer vs. the vanilla Transformer. (Right) The average token similarity of hidden states across layers in Llama (8B) (Dubey et al., 2024) and Mistral (7B) (Jiang et al., 2023).
A.2 Pre-train Dataset
@@ A.3 Training Details
A.4 Validation Loss On Slimpajama

### v2

@@ Abstract
Transformers can capture long-range dependencies using self-attention, allowing tokens to attend to all others directly. However, stacking multiple attention layers leads to attention concentration. One natural way to address this issue is to use cross-layer attention, allowing information from earlier layers to be directly accessible to later layers. However, this approach is computationally expensive. To address this problem, we propose Transformer with residual value (ResFormer) which approximates cross-layer attention through adding a residual connection from the values of the the first layer to all subsequent layers. Based on this method, one variant is the Transformer with single layer value (SVFormer), where all layers share the same value embedding from first layer, reducing the $KV$ cache by nearly 50%. Comprehensive empirical evidence demonstrates that ResFormer mitigates attention concentration problem in deeper layers and enhances representation across most layers, outperforming the vanilla Transformer, DenseFormer, and NeuTRENO in training error as well as downstream tasks. Further visualization results suggest that Resformer alleviates attention sinks through avoiding value-state drains. SVFormer trains significantly faster than the vanilla Transformer and performs better than other methods like GQA and CLA, with performance influenced by sequence length and cumulative learning rate.  [^1] Code is available at https://github.com/Zcchill/Value-Residual-Learning.
@@ 1 Introduction
Existing solutions to alleviate the over-smoothing problem in Transformer include adding extra regularizers (Nguyen et al., 2023, Shi et al., 2022) and optimizing the information flow within the model (Pagliardini et al., 2024). During the era of convolutional neural network architectures, Stochastic Depth (Huang et al., 2016) reduces the likelihood of over-smoothing by randomly dropping layers during training and DenseNet (Huang et al., 2017) mitigates the impact of over-smoothing by allowing each layer to directly access the hidden states of all preceding layers. Recently, DenseFormer (Pagliardini et al., 2024) adopts the idea of DenseNet when training Transformer. Additionally, NeuTRENO (Nguyen et al., 2023) alleviates over-smoothing through incorporating the difference between the value states of the first layer and the current layer to the attention output.
@@
To alleviate the attention concentration from stacking multiple attention layers, an effective method is to use cross-layer attention on information from earlier layers. Given the high computational cost of cross-layer attention, we propose ResFormer as an approximation. ResFormer achieves a similar effect by applying a residual connection between the value states of the current layer and the first layer before the attention operation. Experimental results show that ResFormer mitigates the attention concentration effect in deeper layers, maintaining attention dispersion throughout the sequence, as shown in Fig. 1 (Middle). Further visualization results show that Resformer effectively alleviates attention sinks (Xiao et al., 2023) through avoiding value-state drains (Guo et al., 2024b).
@@
Figure 2: Simplified illustration of the vanilla Transformer, NeuTRENO, DenseFormer, ResFormer, and SVFormer, with only three-layer structures and no operations other than attention. $A^{i}$, $V^{i}$, and $H^{i}$ denote the attention matrix, value states, and attention outputs at the $i$-th layer, respectively. $\oplus$, $\ominus$, and $\otimes$ represent standard matrix addition, subtraction, and multiplication, respectively.
@@ 2.1 Shortcut Connections For Better Information Flow
Related research indicates that for advanced Transformer architectures, although increasing depth continues to yield performance improvements in language modeling tasks, the gains become less significant with further increases (Petty et al., 2024). Furthermore, Zhou et al. (2021) illustrates that a 32-layer ViT underperforms a 24-layer ViT. Depth-Wise Attention (ElNokrashy et al., 2024) allows each query to access the key and value at the same position from previous layers through an attention-like mechanism before the output layer. DenseFormer (Pagliardini et al., 2024) integrates weighted fusion of outputs from all preceding layers after each layer. To explore why increasing depth in Transformers does not yield expected gains, Wang et al. (2022) finds that self-attention acts as a low-pass filter, smoothing token representations in ViTs. Additionally, Shi et al. (2022) investigates over-smoothing from a graph perspective in BERT-based language modeling tasks. NeuTRENO (Nguyen et al., 2023) adds the difference between the value states of the first and current layers to each layer’s attention output and significantly alleviates the over-smoothing problem.
@@ 3.2 Efficient Cross Layer Attention
$\mathbf{U}_{n}$ $=\mathbf{A}_{n}\mathbf{V}_{n}+\frac{1}{n-1}\sum_{i=1}^{n-1}{\mathbf{V}_{i}}.$ (6)
@@ 3.5 SVFormer: Single-layer Value For Half $KV$ Cache
After ResFormer, a natural idea is whether we can remove the value states in each layer and have all layers share the value states from the first layer. We call this method SVFormer. Similar to ResFormer, SVFormer still adopts standard attention in the first layer and obtain the attention output $\mathbf{U_{n}}$ for $n$-th layer where $n\geq 2$ through
@@
Compared to previous methods, SVFormer is the first method that decouple value states from attention. Its main advantage is that it only requires computing and storing the value states for the first layer, saving nearly half of the $KV$ cache during inference. Similar methods like CLA reduce $KV$ cache by sharing both of the key and value states every two layers, but the results in Fig. 4 show that sharing values has less negative impact on performance compared with sharing keys.
@@ 4.1.3 Entropy For Analyzing Attention Concentration Effects
where $\bm{a}_{i}^{\prime}={\bm{a}_{i}}/\Big{(}{\sum_{i=1}^{l}{\bm{a}_{i}}}\Big{)}$ for $i=1,2,\dots,l$ and the higher the entropy $\bm{e}$, the greater the degree of clustering in $\bm{a}$, i.e., attention matrix $\mathbf{A}$ is more likely to focus on several specific tokens.
@@ 4.1.4 Spectral Decomposition For Analyzing Representations
We trained ResFormer and vanilla Transformer with different model size on data with different sequence lengths. In Fig. 5, ResFormer consistently outperforms vanilla Transformer throughout training for different models sizes. Additionally, the results in Fig. 1 (Left) illustrate that ResFormer outperforms DenseFormer and NeuTRENO. Furthermore, integrating ResFormer with NeuTRENO leads to additional performance improvements. We further test the variant of Resformer defined as $\mathbf{U}_{n}=\mathbf{A}_{n}({\mathbf{V}_{n}}+{\bm{\lambda}}{\mathbf{V}_{1}})$. As shown in Fig.6, Resformer can accommodate a wide range of $\bm{\lambda}$ values and the performance improves as $\bm{\lambda}$ increases, achieving the best results at ${\bm{\lambda}}=2$. Regardless of the value of $\bm{\lambda}$, ResFormer consistently outperforms Transformers. It suggests that the success of Resformer lies in the use of $\mathbf{V}_{1}$ and the transformation through $\mathbf{A}_{n}$. The ablation study of different hyperparameters $\bm{\lambda}$ for NeuTRENO, as defined in Equation 8, can be found in the Appendix A.3.     [Uncaptioned image] Figure 6: Ablation study of different $\bm{\lambda}$ for Resformer.
@@ 4.3 Ablation study of residual connection
(a) Token importance.
(b) Norms of value states.
(c) Norms of hidden states.
Figure 10: The token importance, value-state norms, and hidden-state norms of the first token in each sequence across layers in ResFormer vs. the vanilla Transformer.
(a) Transformer.
(b) NeuTRENO.
(c) Resformer.
Figure 11: The distribution of token importance for different models at different layers.
@@ 4.4 Downstream Evaluations
Figure 12: Left: Distribution of eigenvalues for the value states in the first layer of ResFormer and Transformer. Right: Maximum eigenvalue for each layer of ResFormer and Transformer.
@@ 4.5 Visualization of ResFormer
To figure out why ResFormer can achieve better performance on language modeling tasks than vanilla Transformer, we conduct visualization based on the eigenvalue decomposition discussed in Section 4.1.4. After sorting the eigenvalues in descending order, we compute the average eigenvalue for each layer across 1,000 randomly sampled pre-train data examples. The results in Fig. 12 indicate that the value states generated by most layers of the ResFormer exhibit stronger representational capacity compared to those of the vanilla Transformer.
Besides, we analyze the attention concentration effects mentioned in Section 4.1.3 using the same batch of test data. Fig. 1 (Middle) illustrates that the clustering effect of attention increases significantly with the number of layers for the vanilla Transformer, whereas the clustering effect is relatively less pronounced for the ResFormer. We further visualize the attention weights, value-state norms $\|\bm{v}\|_{2}$, and hidden-state norms $\|\bm{h}\|_{2}$ of tokens at different layers and positions, with detailed results in Appendix A.2. Given that attention clustering often occurs on the first token, we primarily show its results in Fig. 10. The results indicate that using ResFormer significantly mitigates attention sinks (Xiao et al., 2023), value-state drains (Guo et al., 2024b) and residual-state peaks (Sun et al., 2024). Guo et al. (2024a) attributes these phenomena to the mutual reinforcement mechanism of model and we suggest that the value shortcut disrupts this mechanism by alleviating value-state drains. Specifically, for tokens lacking semantic information, such as the first token, a large value state magnitude can adversely affect the prediction of future tokens if it is overly attended to, resulting in higher loss.
@@ 5 Conclusion
In this paper, we propose the concept of attention concentration, a problem that arises from stacking multiple attention layers. From the perspective of cross-layer attention, we derive ResFormer, which adds a residual connection between the value states of the current layer and those of the first layer before the attention operation to alleviate attention concentration. Additionally, we introduce SVFormer, based on ResFormer, which reduces the $KV$ cache by nearly half. We conducted comprehensive experiments on the language modeling task to validate the advantages of these two Transformer variants in different scenarios.
@@ A.1 Token Similarity Analysis
Figure 18: The average token similarity of hidden states across layers in Llama and Mistral.
A.2 Attention Concentration Visualization
We visualize the token importance, norms of value states and norms of hidden states for tokens at different position across layers. The results are averaged from 1,000 different sequences so that only the start token is the same and special across all sequences. Fig. 19 (First row) demonstrates that the start token easily attracts massive attention despite lacking semantic information for Transformer and NeuTRENO. For Resformer, the importance of the start token is less than 10 times that of tokens at other positions, indicating that tokens carrying semantic information receive more attention. Moreover, both Transformer and NeuTRENO exhibit significant value-state drains (Guo et al., 2024b) and residual-state peaks (Guo et al., 2024a, Sun et al., 2024) on the start token at certain layers. In contrast, for Resformer, the value state norm of the start token is approximately 0.7 times that of other tokens, while the peak hidden state norm is about three times larger.
Moreover, We conduct an entropy analysis on test examples of varying lengths. The results in Fig.20 indicate that Resformer is more effective in mitigating attention concentration when the sequences are shorter.
Token
importance
Norms of
value states
Norms of
hidden states
(a) Transformer.
(b) NeuTRENO.
(c) Resformer.
Figure 19: Visualization of the token importance, norms of value states and norms of hidden states.
(a) Length: 20 tokens.
(b) Length: 100 tokens.
(c) Length: 500 tokens.
(d) Length: 2,048 tokens.
Figure 20: The average entropy of token importance across layers in ResFormer vs. the vanilla Transformer for sequence with different lengths.
A.3 Ablation study of NeuTRENO
NeuTRENO is sensitive to the choice of hyperparameter $\bm{\lambda}$ which is task-dependent. In the appendix of Nguyen et al. (2023), it is reported that $\lambda$ is set to 0.6 for image classification and segmentation tasks, and 0.4 for language modeling tasks. Fig. 21 indicates that ${\bm{\lambda}}=0.4$ achieves the best results in our training dataset so that we choose ${\bm{\lambda}}=0.4$ for comparison. Besides, we empirically choose ${\bm{\lambda}}=0.2$ for NeuTRENO when combined with Resformer.
Figure 21: Ablation study of different $\lambda$ for NeuTRENO.
A.4 Pre-train Dataset
@@ A.5 Training Details
A.6 Validation Loss On Slimpajama